OpenStudy (anonymous):

If 2x+2y=u and 2xy=v Find x as a function of u,v y as a function of u,v

OpenStudy (anonymous):

Many ways to do this one.

OpenStudy (anonymous):

Square the top equation: \[ (2x+2y)^2=u^2\implies 4x^2+4xy+4y^2=u^2 \]Double the bottom one:\[ 2(2xy) = 2v\implies 4xy = 2v \]Subtract bottom from top: \[ 4x^2+4y^2=u^2-2v \]Divide by \(4\):\[ x^2+y^2=\frac{u^2-2v}{4} \]This is just a circle.

OpenStudy (anonymous):

Hmm, made a small error.

OpenStudy (anonymous):

no x the required is x alone and y alone

OpenStudy (anonymous):

Square the top equation: \[ (2x+2y)^2=u^2\implies 4x^2+8xy+4y^2=u^2 \]Quadruple the bottom one:\[ 4(2xy) = 4v\implies 8xy = 4v \]Subtract bottom from top: \[ 4x^2+4y^2=u^2-4v \]Divide by \(4\):\[ x^2+y^2=\frac{u^2-4v}{4} \]This is just a circle.

OpenStudy (anonymous):

How do you know that it is even possible?

OpenStudy (anonymous):

i do not know

OpenStudy (anonymous):

It's not possible to make \(x\) and \(y\) functions of \(v\) and \(u\).

OpenStudy (anonymous):

Because the relationship isn't that of a function.

OpenStudy (anonymous):

For something as simple as \(y=x^2\), we can make \(y\) a function of \(x\) but \(x\) is not a function of \(y\) because \(y=1\) has \(x=-1\) and \(x=1\) as solutions.

OpenStudy (anonymous):

thank you very much

OpenStudy (ikram002p):

hmm its just a circl lol can u solve it -.-